OpenStudy (anonymous):

I need help on 10 questions and whoever is willing to help could I ask that you walk me through them? I know this is alot to ask and no this is not all my homework Ive done quite alot of it and am stuck on what to do on them. Which of the following represents x^2-25/x^2-5x in simplified form?

OpenStudy (anonymous):

What's the options?

OpenStudy (anonymous):

x-5/x x+5/x 1+5/x 1-5/x

OpenStudy (anonymous):

Soory for asking so much. I just honestly dont know any of this tuff...

OpenStudy (anonymous):

I'd go with the last choice

OpenStudy (anonymous):

nbd

OpenStudy (anonymous):

I chose that one and got it wrong...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

My assignment is to correct all the wrong answers on a quiz to get bonus points (which I need)

OpenStudy (anonymous):

and so the ones I got wrong are ones I dont know anything about. (I got half of them right XD)

OpenStudy (anonymous):

is this algebra 2?

OpenStudy (anonymous):

mo its alg 1

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I tested it out by substituting x for 1 and got a positive number

OpenStudy (anonymous):

Since x in the denominator is negative, I think the negative was factored out of both numerator and denominator, leaving you with a 1+5/x (I'm assuming)

OpenStudy (anonymous):

it sais that x+5/x is wright and Ive tried simplifying it with the exponents and stuff but couldnt do anything...

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

You can square the numbers on the top, which would give you x + 5

OpenStudy (anonymous):

how do you square again? (Im so hopeless XD)

OpenStudy (anonymous):

lol when you multiply a # by itself you get a squared number since 5*5=25, you get five. X^2 is x*x, so you just use the x

OpenStudy (anonymous):

ok but what would you do for the bottom. You cant square 5...

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

I can ask my teacher in monday about it. Want to go on to the next question?

OpenStudy (anonymous):

I have to go, sorry :(

OpenStudy (anonymous):

aw ok

OpenStudy (anonymous):

If you post them, I might be able to do some of them later

OpenStudy (anonymous):

33. What are the exact solutions of x 2 − 3x − 1 = 0? (4 points) (0 pts) x = (0 pts) x = (4 pts) x = (0 pts) x =5/18/13 ue 10/18 0 /4 points 34. Which expression is the simplified form of ? (4 points) (0 pts) 5xy 4 (4 pts) 6xy 4 (0 pts) 5x 2y 8 (0 pts) 6x 2y 8 0 /4 points 35. What is the exact value of the expression ? Simplify completely. (4 points) (0 pts) (0 pts) (4 pts) (0 pts) 0 /4 points 36. What is the product of ? Simplify if possible. (4 points) (0 pts) 294x (4 pts) 42x (0 pts) 12x 2 (0 pts) 42 0 /4 points 37. Blake has a total of 64 DVDs and CD's. If the number of CDs is eight more than three times the number of DVDs, how many DVDs does he have? (4 points) (4 pts) 14 (0 pts) 22 0 /4 points 38. Factor completely: 36x 2 − 49 (4 points) (0 pts) (6x − 1)(6x − 49) (0 pts) (6x + 1)(6x − 49) (4 pts) (6x + 7)(6x − 7) (0 pts) (6x − 7)(6x − 7)5/18/13 ue 11/18 0 /4 points 39. The tickets for a high school basketball game are $2.75 for students and $4.00 for adults. At Tuesday night's game, 132 tickets were sold for a total of $428.00. How many student tickets were sold? (4 points) (0 pts) 78 (4 pts) 80 (0 pts) 50 (0 pts) 52 0 /4 points 40. What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5 (4 points) (4 pts) -1 (0 pts) 5 0 /4 points 41. Choose the system of equations which matches the following graph.

OpenStudy (anonymous):

34. sqrt 108x3y8 35. sqrt 18 + sqrt 48 - sqrt 50 36. 2 sqrt 35x * 3 sqrt 14x

OpenStudy (anonymous):

For the first one in the long list you can use The quadratic formula to solve it for 2 values of x x=\[\frac{ -b \pm \sqrt{b ^{2} -4ac} }{ 2a }\] where a= the number before \[ x ^{2}\] b= number before x c= the plain number in this case a=1 b=-1 c=-1

OpenStudy (anonymous):

Yes that's why i couldn't help with the rest of the questions its very packed together

OpenStudy (anonymous):

Yeah :(